OpenStudy (anonymous):

WRITE IN EXPONETIAL FORM LOG3 81=4

OpenStudy (anonymous):

3^4=81

OpenStudy (anonymous):

essentially log means, what is the number that you raise another number to in order to get this value, here \[\log_{3}81 = 4 \], means, the number you raise 3 to in order to get 81 is 4

OpenStudy (anonymous):

so \[3^4 = 81\] =\[3^2 * 3^2 = 9*9=81\] check!